OpenStudy (anonymous):

y= 1/4x^2 + 3x + 6... how do you complete the square for that?

OpenStudy (amistre64):

factor the 1/4, or mutliply both sides by 4

OpenStudy (anonymous):

4y= x^2 +3x+ 6

OpenStudy (anonymous):

?

OpenStudy (amistre64):

close :) 4y = x^2 +12x +24

OpenStudy (amistre64):

now complete the square as usual and then divide off the 4 if wanted

OpenStudy (anonymous):

oh okay how do you do that?

OpenStudy (amistre64):

do what, complete the square?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

well, lets consider a generic square setup: (x+n)^2 , can you expand that out? to a trinomial setup?

OpenStudy (anonymous):

yes x^2 +2xn+n^2

OpenStudy (amistre64):

perfect, so our format needs to look like this x^2 + 2n x + n^2 ,but they we have x^2 +12 x + 24 by equating like terms we get: x^2 = x^2 2n x = 12 x n^2 = 24; for get this last line at the moment the crucial thing is to determine a suitable value for n: so what is a solution to: 2n = 12 ?

OpenStudy (anonymous):

6

OpenStudy (amistre64):

good, then n^2 = 6^2 = 36; but we cant really go changing the inherent value of the setup, so we would LOVE to add zero to it. it just so happens that 36 - 36 = 0 so lets complete the square for this x^2 +12x + 24 x^2 +12x + 24 + 36 - 36 x^2 +12x + 36 + 24 - 36 (x^2 +12x + 36) + 24 - 36 (x+6)^2 - 12 therefore 4y = (x+6)^2 - 12

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

now, the next step depends on why you would want to complete it to begin with; if its just to place the quadratic in vertex form then divide off the 4 \[y=\frac14(x+6)^2-3\] if its to find the roots .... \[0=(x+6)^2-12\] \[12=(x+6)^2\] \[\pm\sqrt{12}=x+6\] \[-6\pm\sqrt{12}=x\]

OpenStudy (amistre64):

the key to completing a square is that: 2n = middle term setup notice that if we add and subtract the square of half that term (12/2)^2 = 6^2 = 36